How much water should be added to 8ml of 6% saline solution to reduced the convention to 4%
Answer by stanbon(75887)

How much water should be added to 8ml of 6% saline solution to reduced the convention to 4%
-----
Equation:
salt + salt = salt
0.06*8 + 0*x = 0.04(8+x)
--------
6*8 + 0 = 4*8 + 4x
----
4x = 2*8
x = 4 ml (amt of water to add)
------------
Cheers,
Stan H.
------------